An electron is accelerated from rest to a speed 0.8c by applying potential difference. Find it's value.
1
Expert's answer
2020-09-07T08:55:36-0400

solution:

given data:-

rest mass of electron(m0)="9.1\times10^{-31}kg"

charge of electron(q)="1.6\times10^{-19}C"

speed of electrone(v)=0.8c

relation between kinetic energy and potential difrence is


"KE=q\Delta V" .........eq.1

and

"KE=\frac{m_0c^2}{\sqrt1-\frac{v^2}{c^2}}-m_0c^2 ......eq.2"


using equation 1 and 2 we get


potential difference can be written as


"q\Delta V=\frac{m_0c^2}{\sqrt1-\frac{v^2}{c^2}}-m_0c^2"


"\Delta V=\frac{m_0c^2}{q}(\frac{1}{\sqrt{1-\frac{v^2}{c^2}}}-1)"


by putting the value of m0 ,c, v and q


"\Delta V=\frac{9.1\times10^{-31}\times9\times10^{16}}{1.6\times10^{-19}}(\frac{1}{\sqrt{1-\frac{(0.64)c^2}{c^2}}}-1)"


"\Delta V=51.1\times10^4\times\frac{2}{3}"


"=34.06\times10^4"


"\fcolorbox{red}{yellow}{$\Delta V=341\times10^3 volts$}"


therefore potential difference for this electron is 341 KV.